OpenStudy (anonymous):

I'll Give A Medal & Fan You .. ! What Is The Solution Of The Proportion ? 7/9 =m/27

OpenStudy (anonymous):

m=21

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

no problem!

OpenStudy (anonymous):

Can You Help Me With Another One @Raphaelcbm

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

3y-8/12 = y/5

OpenStudy (anonymous):

y=5/21

OpenStudy (anonymous):

Thats Not A Choice

OpenStudy (anonymous):

what are the choices?

OpenStudy (triciaal):

3 y - 8/12 = y/5 will give you y = 5/21 as above however, (3 y-8)/12 = y/5 5(3 y-8)=12 y (15-12)y = 40 y = 40/3 check left side 3(40/3) -8 then divide by 12 = 32/12 reduce = 8/3 right side (40/3)/5 = (40/3)(1/5) divide 40by 5 to reduce = 8/3 please be careful when writing the problem so you can get the proper help